# Magneto-optical properties of Group-IV–vacancy centers in diamond upon hydrostatic pressure

### v1

@@ Abstract
In recent years, the negatively charged group-IV–vacancy defects in diamond, labeled as G4V(--) or G4V centers, have received a great attention in quantum information processing. In this study, we investigate the magneto-optical properties of the G4V centers under high compressive hydrostatic pressures up to 180 GPa. The spin-orbit splitting for the electronic ground and excited states and the hyperfine tensors are calculated by means of plane wave supercell density functional theory as unique fingerprints of these defects. To this end, we developed a theory for calculating the hyperfine tensors when the electronic states are subject to Jahn–Teller effect. We find that the zero-phonon-line energy increases with adding hydrostatic pressures where the coupling strength increases from SiV(--) to PbV(--). On the other hand, the calculated photoionization threshold energies indicate that the operation of PbV(--) based quantum sensor is limited up to 30 GPa of hydrostatic pressure whereas SnV(--), GeV(--) and SiV(--) remain photostable up to 180 GPa of hydrostatic pressure.
@@ I Introduction
The negatively charged group-IV–vacancy (G4V) defects in diamond, i.e., G4V(--) defects or G4V centers, have gained substantial attention in the last decades (see the schematic geometry and electronic structure in Fig. 1). These color centers including silicon-vacancy (SiV), germanium-vacancy (GeV), tin-vacancy (SnV), and lead-vacancy (PbV) defects are promising qubits for quantum communication and sensor applications Vavilov and Gippius (1980); Hepp et al. (2014); Müller et al. (2014); Rogers et al. (2014a); Ekimov et al. (2015); Iwasaki et al. (2015, 2017a, 2017b); Meesala et al. (2018); Ditalia Tchernij et al. (2018); Trusheim et al. (2019); Chen et al. (2019); Trusheim et al. (2020); Görlitz et al. (2020); Krivobok et al. (2020); Rugar et al. (2021); Aghaeimeibodi et al. (2021); Wang et al. (2021); Vindolet et al. (2022); Wang et al. (2024); Bhaskar et al. (2020); Bersin et al. (2024); Knaut et al. (2024). The G4V centers have $S=1/2$ spin state and exhibit $D_{3d}$ inversion symmetry Goss et al. (1996); Gali and Maze (2013); Thiering and Gali (2018) which leads to extraordinary optical stability when compared to other color centers such as nitrogen-vacancy center in diamond Doherty et al. (2013); Gali (2019), which can be utilized in quantum communication technologies Bhaskar et al. (2020); Bersin et al. (2024). The electronic structure of G4V centers provides an advantage to achieve coherent control of spin states using only optical methods, without the need for microwave fields Pingault et al. (2014); Becker et al. (2016); Siyushev et al. (2017); Becker et al. (2018); Weinzetl et al. (2019); Debroux et al. (2021). The ability of microwave-free control of spin states can be practical under extreme conditions such as high pressures where the reduction of the complexity of measurement is a very important issue Vindolet et al. (2022). On the other hand, the effect of pressure on the magneto-optical properties of the G4V centers has not yet been explored in details Vindolet et al. (2022); Harris et al. (2023) which is an inevitable step to determine the potential of these color centers in high-pressure quantum sensor applications.
@@
In this paper, we explore the magneto-optical properties of G4V centers upon compressive hydrostatic pressure by means of density functional theory plane wave supercell calculations. In particular, we focus on the fine level structure of the electronic structure, i.e., the spin-orbit and hyperfine interaction in the respective ground state and excited state. These properties are intertwined with the strong electron-phonon coupling that can be described by the Jahn–Teller theory Thiering and Gali (2018). We developed a theory to accurately compute the hyperfine tensors subject to the Jahn–Teller effect and implemented it to calculate all the critical hyperfine tensor elements of the respective group-IV dopant and the proximate 13C nuclear spins. The photoionization threshold energies are also monitored in our study, along with the shift in the zero-phonon-line energy as a function of compressive hydrostatic pressure. We provide the list of pressure-dependent magneto-optical parameters which indicate the coupling strength of pressure to zero-phonon-line (ZPL) energies, spin-orbit gaps, and hyperfine levels. These parameters can be used to calibrate the actual pressure acting on the color centers. We find that the coupling strength increases when going from lighter to heavier group-IV element in the G4V centers. Our calculations also reveal that the operation of the PbV(--) quantum sensor is limited up to 32 GPa.
@@ II Methods
Figure 1: Schematic diagrams about the basic properties of G4V(--) color centers in diamond. (a) Geometry of the defect. The two vacancies and the interstitial dopant atom (silicon, germanium, tin, or lead) are labeled by dashed spheres and a red sphere, respectively. The critical distances under high $D_{3d}$ symmetry are labeled by $d_{1}$ and $d_{2}$. The dopant atom sits in the inversion center of the diamond. (b) Defect levels in the electronic ground state. The positions of the fully occupied double degenerate $e_{u}$ levels and the partially occupied $e_{g}$ levels vary in the various G4V(--) color centers. In the electronic excited state (not shown) the $e_{g}$ levels are fully occupied whereas a hole is left on the $e_{u}$ level. (c) Electronic structure in the ground and excited states. $\lambda_{g,u}$ are the effective spin-orbit splitting in the $g$ (even parity) ground and $u$ (odd parity) excited states. The effect of the applied magnetic field (significantly smaller than $\lambda_{g,u}$) is also depicted. (d) Adiabatic potential energy surface (APES) of the quadratic Jahn–Teller system with single effective symmetry breaking phonon. $Q_{x,y}$ configuration coordinates represent the effect of the symmetry-breaking degenerate $e_{\text{g}}$ phonon on the APES of which frequency is given by the steepness of the paraboloids. $E_{\text{JT}}$ is the difference in energy between the high symmetry configuration and the distorted configuration. The three equivalent global minima are separated by energy barriers of $\delta_{\text{JT}}$.
@@
We employed density functional theory (DFT) calculations Kohn and Sham (1965); Hohenberg and Kohn (1964) using plane-wave supercell and the projector-augmented-wave (PAW) approach Blöchl (1994); Kresse and Joubert (1999) as implemented in the Vienna Ab-initio Simulation Package (VASP) Kresse and Furthmüller (1996, 1996). The exchange-correlation functional of strongly constrained and appropriately normalized (SCAN) meta-GGA was employed Sun et al. (2015). For the hyperfine tensor calculations, we also list results with HSE06 functional Heyd et al. (2003). The defect structure is simulated within a $4\times 4\times 4$ supercell, which consists of 512 atoms with a lattice constant of 14.21 Å and 14.18 Å in the SCAN and hybrid functional, respectively. $\Gamma$-point was used for sampling the Brillouin-zone. The structures optimized with the criterion of 10-2 eV/Å per atom for the Hellman-Feynman forces and a kinetic energy cutoff of 600 eV was used. The geometry optimization of the electronic excited state is carried out by the $\Delta$SCF method Gali (2009) so the ZPL energies are computed as the total energy difference in the respective electronic states in their global adiabatic potential energy surface minima.
Next, we discuss the spin-orbit coupling and hyperfine tensor calculations which yield the fine electronic structure of G4V(--) defects. We determined the spin-orbit coupling (SOC) with SCAN functional in the non-collinear approach for the G4V centers by using the quantization axis of the spin along the symmetry axis of the defect. The SOC has a small perturbation effect on electronic states, therefore we calculate it in the optimized high $D_{3d}$ symmetry configuration. G4V(--) defects have double degenerate $e_{g\{x,y\}}$ states in the gap that are occupied with three electrons, which can be considered as single hole occupation of the double degenerate $e_{g\{x,y\}}$ states. After applying the SOC on the system, the associated $z$ component of SOC, $\lambda_{0}$, splits the $e_{g\pm}=(\frac{1}{\sqrt{2}})(e_{gx}\pm ie_{gy})$ states. We applied $\Delta$SCF method to either occupy $e_{gx}$ or $e_{gy}$ spinor state (constraint occupation of Kohn-Sham orbitals, e.g., Ref. Gali, 2009). The Hamiltonian of SOC is given by
@@
${\hat{H}_{\text{SOC}}}=-\lambda_{0}\hat{L}_{z}\hat{S}_{z}=$ $\frac{\lambda_{0}}{2}\left[|e^{\upharpoonright}_{g+}\rangle\langle e^{\upharpoonright}_{g+}|+|e^{\downharpoonright}_{g-}\rangle\langle e^{\downharpoonright}_{g-}|\right.$ (1)
@@
$\left.-|e^{\upharpoonright}_{g-}\rangle\langle e^{\upharpoonright}_{g-}|-|e^{\downharpoonright}_{g+}\rangle\langle e^{\downharpoonright}_{g+}|\right]\text{,}$
@@
where the $\hat{L}_{z}$ is the orbital moment operator and $\hat{S}_{z}$ is the $z$ component of the electronic spin. The negative sign of $\lambda_{0}$ occurs because of the hole quasiparticle.
The hyperfine (HF) interaction refers to the coupling between nuclear spins and the electron spin, causing additional splitting of spectral lines in the electron spin resonance spectrum. This provides a unique fingerprint of atomic structure when the spin density has an overlap with the nuclear spins. The hyperfine tensor is expressed as
$A^{(I)}_{ij}=$ $\frac{1}{2S}\int d^{3}r\rho_{s}(r)\gamma_{I}\gamma_{e}\hbar^{2}$
@@
where $\gamma_{e}$ and $\gamma_{I}$ are the electron and nucleus Bohr magneton, respectively. The $\rho_{s}(r)$ is the spin density of the spin state of $S$ at the position $r$ and $r_{i}$ refers to the unit vector of $i={x,y,z}$. The first and second terms represent the Fermi-contact and dipole-dipole terms, respectively. The Fermi-contact term strongly depends on the spin density localized at the nucleus. We note that Eq. (2) is modified within the PAW formalism Blöchl (2000), and the spinpolarization of the core orbitals is considered for accurate Fermi-contact term calculations Szász et al. (2013).
@@
We observed that the dynamic Jahn-Teller (DJT) effect suppresses the orbital moment, at least partially, which is referred to as the Ham-reduction Ham (1965); Bersuker (2006); Bersuker and Polinger (2012). The inherent electronic $\lambda_{0}$ can be read directly from the scalar-relativistic spin-orbit splitting Steiner et al. (2016) of the respective $e_{g}$ or $e_{u}$ Kohn-Sham orbitals. The electronic spin-orbit splitting is suppressed due to DJT by the Ham reduction factor $p$; thus, the observed value in experiments is $\lambda=p\lambda_{0}$. The value of $p$ depends on the vibronic wavefunctions (see the derivation in Ref. Thiering and Gali, 2018), thus the vibronic wavefunctions should be calculated that we implement within DJT effect. DJT also substantially affects the hyperfine tensor parameters that we derive in the Appendix (Sec. B). We show there that static hyperfine tensor calculation in the high symmetry configuration is insufficient to describe the complex physics of the vibronic Jahn–Teller states.
@@
In our study, we explore the Jahn–Teller effect by calculating parameters for an effective single-degenerate-mode model. This model represents motion along a symmetry-breaking direction within the $e_{g}$ irreducible representation, adhering to the methodology outlined in previous works Thiering and Gali (2018). Within the linear Jahn–Teller theory, the $E_{g}\otimes e_{g}$ (electronic ground state) and $E_{u}\otimes e_{g}$ (electronic excited state) APESs have a sombrero shape. The inclusion of quadratic terms disrupts the axial symmetry, resulting in three equivalent minima and saddle points Bersuker and Polinger (2012) as illustrated in Fig. 1(d).
@@
Below, we discuss the ab initio DFT analysis of the DJT effect and the respective magneto-optical properties. To address convergence issues in degenerate high-symmetry configurations and accurately determine $D_{3d}$ symmetry geometries, we adopt a strategy involving half-half fractional occupations in half-filled degenerate orbitals within the spin-minority channel. This method simulates the ensemble of two degenerate states, effectively suppressing JT interactions and relaxing to the $D_{3d}$ configuration. Subsequently shifting to integer occupations enables for relaxation along the $e_{g}$-symmetry direction towards the lowest energy broken-symmetry configuration [BS point in Fig. 1(d)]. This technique enables the identification of the lowest energy point, $Q_{e_{g};\text{BS}}$.
@@
However, this technique sometimes results in divergent energies as $w$ approaches zero from positive and negative directions due to the inability of the exchange-correlation functional to yield identical energies for degenerate densities Bruyndonckx et al. (1997). We address this issue by applying a rigid vertical shift to one of the branches, ensuring identical energies at $w=0$.
Assuming only linear Jahn–Teller interaction, to simplify the estimation of Jahn–Teller relaxation energy $E_{\mathrm{JT}}$ at different pressures and to reduce the computational complexity, we develop a novel practical computational procedure that avoids issues related to degeneracy of electronic states. Since the slice cut of sombrero APES is described by $U(Q_{e_{g};x})=\omega^{2}Q_{e_{g};x}^{2}/2-VQ_{e_{g};x}$, where $V$ is vibronic coupling constant, energy minimum is at $\Delta Q_{e_{g};\text{BS}}=V/\omega^{2}$, yielding Jahn–Teller relaxation energy $E_{\mathrm{JT}}=V^{2}/2\omega^{2}$ Bersuker and Polinger (2012). However, by using fractional occupations described above we suppress Jahn–Teller interactions and potential energy surface attains the harmonic form $U_{\text{ad}}(Q_{gx})=\omega^{2}Q_{e_{g};x}^{2}/2$. Plugging in the coordinate of Jahn–Teller minimum into harmonic potential we obtain $U_{\text{ad}}(\Delta Q_{e_{g};\text{BS}})=V^{2}/2\omega^{2}\equiv E_{\mathrm{JT}}$. This allows to estimate $E_{\mathrm{JT}}$ from high-symmetry and broken-symmetry geometries and one single-point calculation with factional occupations at broken-symmetry geometry.

### v2

@@ Abstract
In recent years, the negatively charged group-IV–vacancy defects in diamond, labeled as G4V(--) or G4V centers, have attracted significant attention in quantum information processing. In this study, we investigate the magneto-optical properties of G4V centers under high compressive hydrostatic pressures up to 180 GPa. The spin-orbit splitting of the electronic ground and excited states, as well as the hyperfine tensors, are calculated using plane-wave supercell density functional theory, providing distinctive fingerprints that uniquely characterize these defects. To this end, we develop a theory for calculating the hyperfine tensors when the electronic states are subject to the Jahn–Teller effect. We find that the zero-phonon-line energy increases with hydrostatic pressure, with the deformation potential increasing from SiV(--) to PbV(--). On the other hand, our calculated photoionization threshold energies indicate that PbV(--)-based quantum sensors can operate only up to 32 GPa, whereas SnV(--), GeV(--), and SiV(--) remain photostable up to 180 GPa. We also find that the spin-orbit splitting increases in both the electronic ground and excited states with increasing pressure. The optical transitions associated with the hyperfine fine structure of the dopant atoms are interpreted using our theoretical framework, which reproduces existing experimental data at zero strain. We show that the hyperfine levels are weakly dependent on magnetic field, and increasing pressure leads to optical transitions at longer wavelengths. Finally, we estimate the spin coherence times of the G4V centers under increasing hydrostatic pressure across different temperature regimes.
@@ I Introduction
The negatively charged group-IV–vacancy G4V(--) defects in diamond, commonly referred to as G4V centers, have attracted significant attention over the past decades due to their promising optical properties (see Fig. 1 for a schematic representation of their geometry and electronic structure). These color centers, including silicon-vacancy (SiV), germanium-vacancy (GeV), tin-vacancy (SnV), and lead-vacancy (PbV) defects are promising qubits for quantum communication and sensor applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. The G4V centers have a $S=1/2$ spin state and exhibit $D_{3d}$ inversion symmetry [24, 25, 26], leading to outstanding optical stability compared to other color centers, such as the nitrogen-vacancy center in diamond [27, 28], which can be utilized in quantum communication technologies [21, 22]. The electronic structure of G4V centers provides an advantage in achieving coherent control of spin states using only optical methods without the need for microwave fields [29, 30, 31, 32, 33, 34]. The ability of microwave-free control of spin states can be practical under extreme conditions such as high pressures, where reducing the complexity of measurement is a very important issue [19]. On the other hand, the effect of pressure on the magneto-optical properties of the G4V centers has not yet been explored in detail [19, 35], which is an inevitable step in determining the potential of these color centers in high-pressure quantum sensor applications.
@@
In this paper, we explore the magneto-optical properties of G4V centers upon compressive hydrostatic pressure using density functional theory (DFT) plane wave supercell calculations. In particular, we focus on the fine-level structure of the electronic system, i.e., the spin-orbit and hyperfine interaction in the respective ground and excited states. These properties are intertwined with the strong electron-phonon coupling that can be described by the Jahn–Teller theory [26]. We developed a theory to accurately compute the hyperfine tensors subject to the Jahn–Teller effect. We implemented it to calculate all the critical hyperfine tensor elements of the respective group-IV dopant and the proximate 13C nuclear spins. The photoionization threshold energies are also monitored in our study, along with the shift in the zero-phonon-line energy as a function of compressive hydrostatic pressure. We provide the list of pressure-dependent magneto-optical parameters that indicate the coupling strength of pressure to zero-phonon-line (ZPL) energies, spin-orbit gaps, and hyperfine levels. These parameters can be used to calibrate the actual pressure acting on the color centers. We find that the coupling strength increases when going from lighter to heavier group-IV elements in the G4V centers. Our calculations also reveal that the operation of the PbV(--) quantum sensor is limited up to 32 GPa.
@@ II Methods
Figure 1: Schematic diagrams illustrating the fundamental properties of G4V(--) color centers in diamond. (a) Geometry of the defect. Dashed spheres represent the two vacancies, and the interstitial dopant atom (silicon, germanium, tin, or lead) is shown as a red sphere. The critical distances under high $D_{3d}$ symmetry are labeled as $d_{1}$ and $d_{2}$. The dopant atom is positioned at the inversion center of the diamond. (b) Schematic single-particle defect levels in the electronic ground state. The positions of the fully occupied double degenerate $e_{u}$ levels and the partially occupied $e_{g}$ levels vary in the various G4V(--) color centers. In the electronic excited state (not shown), the $e_{g}$ levels are fully occupied, whereas a hole is left on the $e_{u}$ level. (c) Electronic structure in the ground and excited states, including fine and hyperfine interactions. The hyperfine structure originates from the coupling with the impurity atom positioned at the inversion symmetry point. The fine and hyperfine states are labeled according to the double-group representations of the $D_{3d}$ point group, following the notation in Ref. [36]. $\lambda_{g,u}$ are the effective spin-orbit splitting in the $g$ (even parity) ground and $u$ (odd parity) excited states. Transitions between two Kramers states occur with orbital relaxation at rates $\gamma_{\pm}$. The hyperfine coupling parameters $A_{\parallel,\{g,u\}}$ and $A_{2,\{g,u\}}$ lift the degeneracy of the fine-structure states, while $A_{\perp,\{g,u\}}$ and $A_{1,\{g,u\}}$ facilitate the mixing of spin-orbit states. (d) Adiabatic potential energy surface (APES) of the quadratic Jahn-Teller system along the ionic degrees of freedom corresponding to a single degenerate $e_{g}$-symmetry mode, described by the configuration coordinates $Q_{x,y}$. The points of BS, HS, and S represent the lowest energy broken-symmetry, the highest energy of highest symmetry and saddle point configurations, respectively. $E_{\text{JT}}$ is the difference in energy between the high symmetry configuration and the distorted configuration. The three equivalent global minima are separated by energy barriers of $\delta_{\text{JT}}$.
@@
We performed density functional theory (DFT) calculations [37, 38] using plane-wave supercell and the projector-augmented-wave (PAW) approach [39, 40] as implemented in the Vienna Ab-initio Simulation Package (VASP) [41, 42]. For the exchange-correlation functional, we employed the strongly constrained and appropriately normed (SCAN) meta-GGA functional [43], which has demonstrated accurate performance in describing structural and electronic properties of color centers in diamond [44]. For the hyperfine tensor calculations, we additionally provide results obtained using the hybrid HSE06 functional [45]. The defect structure was simulated within a $4\times 4\times 4$ supercell consisting of 512 atoms, with lattice constants of 14.21 Å and 14.18 Å obtained using the SCAN and hybrid functionals, respectively. $\Gamma$-point was used to sample the Brillouin zone. The structures were optimized using a convergence criterion of 10-2 eV/Å per atom for the Hellman-Feynman forces, with a kinetic energy cutoff of 600 eV.
The G4V(--) defects feature double-degenerate $|e_{u}\rangle$ and $|e_{g}\rangle$ localized single-particle levels, corresponding to irreducible representations of the $D_{3d}$ symmetry group [see Fig. 1(b)]. These levels lie within the bandgap and are occupied by seven electrons. The ground-state configuration can be described as a single-hole occupation of the $|e_{g}\rangle$ orbitals, giving rise to the spin-doublet ${}^{2}\!E_{g}$ state [see Fig. 1(c)]. The excited state is formed by promoting an electron from an orbital with $e_{u}$ symmetry to the $|e_{g}\rangle$ orbital, resulting in the ${}^{2}\!E_{u}$ state. Geometry optimization of the electronic excited state was conducted using the $\Delta$SCF method [46], with the Kohn–Sham orbital occupation constrained to achieve the ${}^{2}\!E_{u}$ configuration. ZPL energies were computed as the total energy difference between the respective electronic states at their global minima on the adiabatic potential energy surfaces. The orbital degeneracy in both the ground and excited states induces Jahn–Teller coupling, which in adiabatic DFT calculations manifests as symmetry-breaking distortions along the nuclear degrees of freedom associated with the $|e_{g}\rangle$ symmetry [26].
Next, we address the calculation of spin-orbit coupling (SOC) and hyperfine tensors, which provide insights into the fine electronic structure of G4V(--) defects. We determined the SOC using the SCAN functional within a non-collinear framework of DFT, with the quantization axis of the spin aligned along the symmetry axis of the defect. The SOC has a small perturbation effect on electronic states; therefore, it was calculated using the optimized high-symmetry $D_{3d}$ configuration. After applying SOC to the system, the associated $z$-component of the SOC, $\lambda_{0}$, splits the ${}^{2}\!E_{g}$ and ${}^{2}\!E_{u}$ states into two degenerate doublets. Within the hole notation, these doublets are defined as $|E_{3/2,\{g,u\}}\rangle=\{|e_{\{g,u\}+}^{\uparrow}\rangle,|e_{\{g,u\}-}^{\downarrow}\rangle\}$ and $|E_{1/2,\{g,u\}}\rangle=\{|e_{\{g,u\}-}^{\uparrow}\rangle,|e_{\{g,u\}+}^{\downarrow}\rangle\}$ [see Fig. 1(c)], where, for example, $|e_{g+}^{\uparrow}\rangle\equiv|e_{g+}\rangle|{\uparrow}\rangle$ denotes the spinor orbitals. Under the operations of the $D_{3d}$ symmetry group, the $|e_{u\pm}\rangle$ orbitals are chosen to transform as the spherical harmonics $Y_{1,\pm 1}\sim$$\mp(x\pm iy$)” of the Condon-Shortley phase convention [47, 36], while the $|e_{g\pm}\rangle$ orbitals transform equivalently to angular momentum raising and lowering operators (axial vectors) $\hat{L}_{\pm}=\hat{L}_{x}\pm i\hat{L}_{y}$. The Hamiltonian for SOC is given by
${\hat{H}_{\text{SOC}}}$ $=-\lambda_{\{g,u\}}\hat{\sigma}_{z}\hat{S}_{z}$
@@
$=\frac{\lambda_{\{g,u\}}}{2}\left[\bigl{|}e^{\uparrow}_{\{g,u\}+}\bigr{\rangle}\bigr{\langle}e^{\uparrow}_{\{g,u\}+}\bigl{|}+\bigl{|}e^{\downarrow}_{\{g,u\}-}\bigr{\rangle}\bigl{\langle}e^{\downarrow}_{\{g,u\}-}\bigr{|}\right.$
@@
$\quad\left.-\bigl{|}e^{\uparrow}_{\{g,u\}-}\bigr{\rangle}\bigl{\langle}e^{\uparrow}_{\{g,u\}-}\bigr{|}-\bigl{|}e^{\downarrow}_{\{g,u\}+}\bigr{\rangle}\bigl{\langle}e^{\downarrow}_{\{g,u\}+}\bigr{|}\right]\text{,}$ (1)
@@
where the $\hat{\sigma}_{z}=\hat{L}_{z}$ represents the orbital angular momentum operator: $\hat{\sigma}_{z}|e_{\pm}\rangle=\pm|e_{\pm}\rangle$, and $\hat{S}_{z}$ is the $z$ component of the electronic spin. The negative sign of $\lambda_{\{g,u\}}$ occurs because of the hole quasiparticle.
The hyperfine (HF) interaction arises from the coupling between nuclear and electron spins, contributing additional splitting in the electron spin resonance spectrum. This interaction is particularly significant when the electron spin density overlaps with the nuclear positions, providing a characteristic spectroscopic fingerprint of the atomic structure. In systems with orbital degeneracy, the HF interaction also couples to degenerate orbital components, making interaction dependent on the vibronic states of the Jahn–Teller system.
For the hyperfine interaction with the central Group-IV ion, the HF Hamiltonian is invariant under the $D_{3d}$ double group symmetry operations. Ensuring that the HF Hamiltonian transforms as a scalar under this symmetry results in the canonical form (for a group theoretical derivation, see Appendix C):
$\hat{H}_{\text{HF}}$ $=\left[\frac{1}{2}A_{\perp\{g,u\}}(\hat{S}_{+}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{+})+2A_{\parallel\{g,u\}}\hat{S}_{z}\hat{I}_{z}\right]\hat{\mathbb{I}}$
$-A_{1\{g,u\}}\left[\left(\hat{S}_{z}\hat{I}_{+}+\hat{S}_{+}\hat{I}_{z}\right)\!\hat{\sigma}_{+}\!+\!\left(\hat{S}_{z}\hat{I}_{-}+\hat{S}_{-}\hat{I}_{z}\right)\!\hat{\sigma}_{-}\right]$
$+\frac{A_{2\{g,u\}}}{2}\left[\hat{S}_{-}\hat{I}_{-}\hat{\sigma}_{+}+\hat{S}_{+}\hat{I}_{+}\hat{\sigma}_{-}\right]\text{.}$ (2)
Here, $\hat{S}_{i}$ and $\hat{I}_{i}$ denote the spin projection and spin raising or lowering operators acting on the electronic and nuclear spin degrees of freedom, respectively. The operators $\mathbb{I}$ and $\sigma_{\pm}=|e_{\{g,u\}\pm}\rangle\langle e_{\{g,u\}\mp}|$ represent the identity operator and the orbital angular momentum raising and lowering operators, respectively, acting within the two-dimensional subspace of degenerate orbitals. Here, we note that while $\hat{\sigma}_{z}$ is equivalent with $\hat{L}_{z}$ the ladder operators $\hat{\sigma}_{\pm}$’s are not the same as $\hat{L}_{\pm}$. When one looks into Eq. (2) more closely angular momentum conservation appears to be violated. However, one has to take into consideration that $\hat{\sigma}_{\mp}$ flips the ”real” angular momentum twice: $\hat{\sigma}_{\mp}|e_{\{g,u\}\pm}\rangle\sim\hat{L}_{\mp}^{2}|e_{\{g,u\}\pm}\rangle\sim\hat{L}_{\mp}|a_{\{1g,2u\}}\rangle\sim|e_{\{g,u\}\mp}\rangle$ because one has to flip though $|a_{\{1g,2u\}}\rangle$ orbitals first. However, $|a_{\{1g,2u\}}\rangle$ lie deep in the valence band [48, 49] thus we omit $\hat{L}_{\pm}$s from discussion from now on. Finally, we note that Eq. (2) neglects the orbital hyperfine interaction, which is justified because the spin density does not build up from the atomic-like orbitals of the dopant atom.
Figure 1(c) illustrates the hyperfine energy level structure arising from the interaction with the impurity atoms in G4V defects, with the corresponding states labeled according to their irreducible representations. The hyperfine interaction splits the $E_{3/2,\{g,u\}}$ levels into two doublets of $E_{\{g,u\}}$ symmetry. In contrast, the upper $E_{1/2,\{g,u\}}$ branch is divided into a doublet of $E_{\{g,u\}}$ symmetry and two singlets of $A_{1,\{g,u\}}$ and $A_{2,\{g,u\}}$ symmetries.
The first term in Eq. (2) is isotropic with respect to the orbital degrees of freedom and remains unaffected by the ionic motion associated with the Jahn–Teller system. As such, it can be regarded as a “static” contribution to the HF interaction, unaffected by the ionic motion. Within this term, $A_{\parallel,\{g,u\}}$ induces splitting between spin–orbit-coupled (SOC) branches, while $A_{\perp,\{g,u\}}$ facilitates coupling between SOC eigenstates, as illustrated by the red dashed arrow in Fig. 1(c).
In contrast, the last two terms of Eq. (2) describe interactions that couple degenerate components of the orbital states via the $\hat{\sigma}_{\pm}$ operators. These contributions depend explicitly on the orbital composition and are sensitive to the vibronic state of the defect, reflecting a dynamical HF coupling. The parameter $A_{1,\{g,u\}}$ enables the mixing of $E_{\{g,u\}}$ states, as indicated by green dashed arrows in Fig. 1(c), while $A_{2,\{g,u\}}$ induces splitting between the $A_{1,\{g,u\}}$ and $A_{2,\{g,u\}}$ levels. In this study, we compute the values of $A_{1}$ and $A_{2}$ for the lowest vibronic state using the reduction factors proposed by Ham [50].
For a selected spin density $\rho_{\alpha}$, the hyperfine (HF) tensor parameters can be computed using the expression:
$\mathcal{A}_{ij}^{(\alpha)}=$ $\frac{\mu_{0}\hbar^{2}}{4\pi}\frac{\gamma_{I}\gamma_{e}}{2\langle\hat{S}_{z}\rangle}\int d^{3}r\rho_{\alpha}(r)$
@@
where $\gamma_{e}$ and $\gamma_{I}$ are the gyromagnetic ratios of the electron and nucleus, respectively; $\rho_{\alpha}(r)$ is the spin density at position $r$, $\langle\hat{S}_{z}\rangle$ is the expectation value of the $z$ component of the total electronic spin, and $r_{i}$ denotes the $i=x,y,z$ components of the position vector relative to the nucleus. The first term represents the Fermi-contact interaction, which strongly depends on the spin density localized at the nucleus, while the second term corresponds to the dipole–dipole interaction. In VASP, the HF tensor expression in Eq. (3) is modified within the projector-augmented wave (PAW) formalism [51], and the spin polarization of the core orbitals is included to ensure accurate calculation of the Fermi-contact term [52].
The HF parameters in Eq. (2) can be calculated using the real-valued representation of $e$-orbitals. The parameter values, derived for a coordinate system with the $z$-axis aligned along the symmetry axis of the $D_{3d}$ group, are as follows [see Eq. (15) in Appendix C and Ref. 53]:
$A_{\parallel}=\tfrac{1}{2}\mathcal{A}_{zz}^{x},\quad A_{\perp}=\frac{1}{2}(\mathcal{A}_{xx}^{x}+\mathcal{A}_{yy}^{x})$
$A_{1}=q\mathcal{A}_{xz}^{x},\quad A_{2}=q\left(\mathcal{A}_{yy}^{x}-\mathcal{A}_{xx}^{x}\right)\text{,}$ (4)
where the subscript $x$ indicates that the HF tensor [Eq. (3)] was computed for the hole spin density constrained to the $|e_{x}\rangle$ orbital of $C_{3v}$. The factor $q$ represents vibronic reduction factors [50, 54], which account for the quenching of the HF interaction caused by the dynamic Jahn–Teller (DJT) effect in the ground vibronic state.
When the HF interaction involves nearby 13C ions, the overall symmetry is reduced to $C_{2h}$, complicating the calculation of HF parameters. In Appendix D we outline the procedure for estimating the HF parameters under this reduced symmetry.
@@
The DJT effect suppresses the orbital moment, at least partially, which is referred to as the Ham-reduction [50, 55, 56]. The inherent electronic $\lambda_{0}$ can be read directly from the scalar-relativistic spin-orbit splitting [57] of the respective $e_{g}$ or $e_{u}$ Kohn-Sham orbitals. The electronic spin-orbit splitting is suppressed due to DJT by the Ham reduction factor $p$; thus, the observed value in experiments is $\lambda=p\lambda_{0}$. The value of $p$ depends on the vibronic states, which we compute by explicitly incorporating the DJT effect as implemented in this study (see the derivation in Ref. 26).
@@
In our study, we explore the Jahn–Teller effect by calculating parameters for an effective single-degenerate-mode model. This model represents motion along a symmetry-breaking direction within the $e_{g}$ irreducible representation, adhering to the methodology outlined in previous works [26]. Within the linear Jahn–Teller theory, the $E_{g}\otimes e_{g}$ (electronic ground state) and $E_{u}\otimes e_{g}$ (electronic excited state) APESs have a sombrero shape. The inclusion of quadratic terms disrupts the axial symmetry, resulting in three equivalent minima and saddle points [56], as illustrated in Fig. 1(d).
@@
Below, we discuss the ab initio DFT analysis of the DJT effect and the respective magneto-optical properties. To address convergence issues in degenerate high-symmetry configurations and accurately determine $D_{3d}$ symmetry geometries, we adopt a strategy involving half-half fractional occupations in half-filled degenerate orbitals within the spin-minority channel. This method simulates the ensemble of two degenerate states, effectively suppressing JT interactions and relaxing to the $D_{3d}$ configuration [HS point in Fig. 1(d)]. Subsequently, shifting to integer occupations enables relaxation along the $e_{g}$-symmetry direction towards the lowest energy broken-symmetry configuration [BS point in Fig. 1(d)]. This technique enables the identification of the lowest energy point, $Q_{e_{g};\text{BS}}$.
@@
However, this technique sometimes results in divergent energies as $w$ approaches zero from positive and negative directions due to the inability of the exchange-correlation functional to yield identical energies for degenerate densities [58]. We address this issue by applying a rigid vertical shift to one of the branches, ensuring identical energies at $w=0$.
Assuming only linear Jahn–Teller interaction, to simplify the estimation of Jahn–Teller relaxation energy $E_{\mathrm{JT}}$ at different pressures and to reduce the computational complexity, we develop a novel practical computational procedure that avoids issues related to the degeneracy of electronic states. Since the slice cut of sombrero APES is described by $U(Q_{e_{g};x})=\omega^{2}Q_{e_{g};x}^{2}/2-VQ_{e_{g};x}$, where $V$ is vibronic coupling constant, energy minimum is at $\Delta Q_{e_{g};\text{BS}}=V/\omega^{2}$, yielding Jahn–Teller relaxation energy $E_{\mathrm{JT}}=V^{2}/2\omega^{2}$ [56]. However, using fractional occupations described above, we suppress Jahn–Teller interactions, and the potential energy surface attains the harmonic form $U_{\text{ad}}(Q_{gx})=\omega^{2}Q_{e_{g};x}^{2}/2$. Plugging in the coordinate of the Jahn–Teller minimum into the harmonic potential we obtain $U_{\text{ad}}(\Delta Q_{e_{g};\text{BS}})=V^{2}/2\omega^{2}\equiv E_{\mathrm{JT}}$. This approach allows us to estimate $E_{\mathrm{JT}}$ from high-symmetry and broken-symmetry geometries and one single-point calculation with factional occupations at broken-symmetry geometry.